OpenStudy (anonymous):

How would i write this equation? Harry has a jar that contains 75% of apple juice and the rest water. The bottle has 1 liter of water. Part A: Write an equation in one variable that can be used to find the total number of liters of apple juice and water in the jar. Define the variable used in the equation. (5 points)

OpenStudy (anonymous):

I know there is 3 liters of apple juice , im just stuck on how to write this equation

OpenStudy (anonymous):

well 3/4 of the jar is juice then 1/4 of the jar is water and there is 1 liter of water then there has to be 4 liters so juice would have to be 3 liters 3 liters of juice is in the jar

OpenStudy (anonymous):

thats not what i was asking , i said i knew there was 3 liters , but HOW to write the equation?

OpenStudy (rock_mit182):

percentage of water = 100 -75 = 25 1 liter represents the 25% x liter represtns 75% =3 A ( amount of litetrs per mixture) 0.75A + 0.25A = 100%

OpenStudy (anonymous):

Wow , why was i thinking it was alot harder , thank you !

OpenStudy (rock_mit182):

A ( percentage of liter per mixture) would be better describing the situation

OpenStudy (rock_mit182):

You're welcome, anyway there are multiple ways to aproach this problem